# TASAR: Transfer-based Attack on Skeletal Action Recognition

### v1

@@ A Post-train Bayesian Perspective on Attack
$\underset{\theta^{\prime}}{\max}\,\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)}p\left(y\mid\mathbf{x},\theta,\theta^{\prime}\right)$
$\approx\underset{\theta_{k}^{\prime}}{\max}\,\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\mathbf{x},\theta,\theta_{k}^{\prime}\right),\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)$ (3)
where $K$ is the number of appended models. Correspondingly, Eq. 2 can be approximately solved by performing attacks on the ensemble of tiny appended models:
$\underset{\|\delta\|_{p}\leq\epsilon}{\arg\min}\,\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\mathbf{x}+\delta,\theta,\theta_{k}^{\prime}\right),\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)$ (4)
@@
Post-train Dual Bayesian Motion Attack
@@
In our preliminary experiments, we found that a naive application of post-train Bayesian attack (Eq. 4) already surpassed the adversarial transfer performance of existing S-HAR attacks, which demonstrates the effectiveness of smoothening the loss surface of surrogates. However, its performance remains slightly inferior to the Bayesian attack via re-training a Bayesian surrogate (Li et al. 2023)(Eq. 2). This performance gap is understandable, as we avoid the prohibitively slow process of sampling the original posterior distribution $\theta\sim p(\theta\mid\mathcal{D})$ by using a tiny Bayesian component for post-training instead. To further eliminate the trade-off between attack strength and efficiency, we propose a novel post-train dual Bayesian optimization for smoothed posterior sampling, to sample the appended models with high smoothness for better transferability (Fig. 2). Moreover, unlike previous transfer-based attacks that assume each frame is independent and ignore the temporal dependency between sequences, we integrate motion dynamics information into the Bayesian attack gradient to disrupt the spatial-temporal coherence of S-HAR models. We name our method Post-train Dual Bayesian Motion Attack.
Post-train Dual Bayesian Optimization.
@@
This motivation is based on the view that models sampled from a smooth posterior, along with the optimal approximate posterior estimating this smooth posterior, have better smoothness (Nguyen et al. 2024). To this end, we aim for proposing a smooth posterior for learning post-train BNNs, hence possibly possessing higher adversarial transferability. Specifically, inspired by the observation that randomized weights often achieve smoothed weights update (Izmailov et al. 2018; Dziugaite and Roy 2017; Jin et al. 2023), we add Gaussian noise to smooth the appended network weights. This is achieved by a new post-train dual Bayesian optimization:
@@
$\underset{\theta^{\prime}}{\max}\,\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)}\mathbb{E}_{\Delta\theta^{\prime}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right)}p\left(y\mid\mathbf{x},\theta,\theta^{\prime}+\Delta\theta^{\prime}\right)$ (5)
@@
For any appended model sampled from the posterior, Eq. 5 ensures that the neighborhood around the model parameters has uniformly low loss. We further use dual Monte Carlo sampling to approximate Eq. 5:
$\min_{{\theta^{\prime}_{k}}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)}\frac{1}{MK}\sum_{k=1}^{K}\sum_{m=1}^{M}L\left(\mathbf{x},y,\theta,{\theta}_{k}^{\prime}+\Delta\theta^{\prime}_{km}\right)\text{, }$
$\Delta\theta^{\prime}_{km}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right)$ (6)
where $L$ is the classification loss. Considering dual MCMC samplings computationally intensive, we instead consider the worst-case parameters from the posterior, followed by (Li et al. 2023). Hence Post-train Dual Bayesian Optimization. can be equivalent to a min-max optimization problem, written as:
@@
$\min_{\theta^{\prime}_{k}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)}\max_{\Delta\theta^{\prime}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right)}\frac{1}{K}\sum_{k=1}^{K}L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}+\Delta\theta^{\prime}\right)\text{, }$
$p(\Delta\theta^{\prime})\geq\xi$ (7)
The confidence region of the Gaussian posterior is regulated by $\xi$. We discuss the sensitivity to $\xi$ in the Appendix B. The entanglement between $\theta^{\prime}$ and $\Delta\theta^{\prime}$ complicates gradient updating. To simplify this issue, we utilize Taylor expansion at $\theta^{\prime}$ to decompose the two components:
@@
$\min_{\theta^{\prime}_{k}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)}\max_{\Delta\theta^{\prime}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right)}\frac{1}{K}\sum_{k=1}^{K}[L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}\right)$
@@
$+\nabla_{\theta^{\prime}_{k}}L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}\right)^{T}\Delta\theta^{\prime}],p(\Delta\theta^{\prime})\geq\xi$ (8)
@@
Since $\Delta\theta^{\prime}$ is sampled from a zero-mean isotropic Gaussian distribution, the inner maximization can be solved analytically. We introduce the inference details and mathematical deduction in Appendix A. As shown in Fig. 2, the loss landscape optimized by post-train Dual Bayesian is significantly smoother than vanilla post-train Bayesian.
Temporal Motion Gradient in Bayesian Attack.
@@
Post-train Dual Bayesian Motion Attack can be performed with gradient-based methods such as FGSM (Goodfellow, Shlens, and Szegedy 2014):
@@
$\tilde{\mathbf{x}}=\mathbf{x}+\alpha\cdot\operatorname{sign}(\sum_{k=1}^{K}\sum_{m=1}^{M}\nabla L\left(\mathbf{x},y,\theta,{\theta}_{k}^{\prime}+\Delta\theta^{\prime}_{km}\right))$ (9)
@@
where $\alpha$ is the step size. For notational simplicity, we notate the classification loss $L\left(\mathbf{x},y,\theta,{\theta}_{k}^{\prime}+\Delta\theta^{\prime}_{km}\right)$ as $L\left(\mathbf{x}\right)$. Assume a motion with $t$ frames $\mathbf{x}=\left[x_{1},x_{2},\cdots,x_{t}\right]$, this attack gradient consists of a set of partial derivatives over all frames $\nabla L(\mathbf{x})=\left[\frac{\partial L(\mathbf{x})}{\partial x_{1}},\frac{\partial L(\mathbf{x})}{\partial x_{2}},\cdots,\frac{\partial L(\mathbf{x})}{\partial x_{t}}\right]$. The partial derivative $\frac{\partial L(\mathbf{x})}{\partial x_{t}}$ assumes each frame is independent, ignoring the dependency between frames over time. This assumption is reasonable for attacks on static data such as PGD (Madry et al. 2017) while infeasible for skeletal motion attacks. In skeletal motion, most S-HAR models learn the spatial-temporal features (Yan, Xiong, and Lin 2018), hence considering motion dynamics in the computing of attack gradient can disrupt the spatial-temporal coherence of these features, leading to more general transferability. To fully represent the motion dynamics, first-order (velocity) gradient $(\nabla L\left(\mathbf{x}\right))_{d1}$ and second-order (acceleration) gradient information $(\nabla L\left(\mathbf{x}\right))_{d2}$ should also be considered. To this end, we augment the original position gradient with the motion gradient, then Eq. 4 becomes:
@@
$\tilde{\mathbf{x}}=\mathbf{x}+\alpha\cdot\operatorname{sign}(\sum_{k=1}^{K}\sum_{m=1}^{M}\nabla L_{dyn}\left(\mathbf{x}\right))$ (10)
$\nabla L_{dyn}\left(\mathbf{x}\right)=\sum_{n=0}^{2}w_{n}(\nabla L\left(\mathbf{x}\right))_{dn},\sum_{n=0}^{2}w_{n}=1$ (11)
@@
where $(\nabla L\left(\mathbf{x}\right))_{d0}=\nabla L(\mathbf{x})$. Motion gradient can be computed by explicit modeling (Xia et al. 2015) or implicit learning (Tang et al. 2022). Given that implicit learning requires training an additional data-driven model to learn the motion manifold, which increases computational overhead, we opt for explicit modeling. Inspired by (Lu et al. 2023), we employ time-varying autoregressive models (TV-AR)(Bringmann et al. 2017) because TV-AR can effectively estimate the dynamics of skeleton sequences by modeling the temporary non-stationary signals (Xia et al. 2015). We first use first-order TV-AR($f_{d1}$) and second-order TV-AR($f_{d2}$) to model human motions respectively:
$f_{d1}:\tilde{x}^{i}_{t}$ $=A_{t}\cdot\tilde{x}^{i}_{t-1}+B_{t}+\gamma_{t}$ (12)
$f_{d2}:\tilde{x}^{i}_{t}$ $=C_{t}\cdot\tilde{x}^{i}_{t-1}+D_{t}\cdot\tilde{x}^{i}_{t-2}+E_{t}+\gamma_{t}$ (13)
@@
where the model parameters $\beta_{t}^{1}=\left[A_{t},B_{t}\right]$ and $\beta_{t}^{2}=\left[C_{t},D_{t},E_{t}\right]$ are all time-varying parameters and determined by data-fitting. $\gamma_{t}$ is a time-dependent white noise representing the dynamics of stochasticity. Using Eq. 12, the first-order motion gradient can be derived as:
$\left(\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t-1}}\right)_{d1}=\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t-1}}+\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t}}\cdot A_{t}$ (14)
@@
Similarly, second-order dynamics can be expressed as below by using Eq. 13:
$\left(\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t-2}}\right)_{d2}=\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t-2}}+\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t-1}}\cdot C_{t-1}$
$+\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t}}\cdot\left(D_{t}+C_{t}\cdot C_{t-1}\right)$ (15)
@@
Experiments
RobustBenchHAR Settings
To our best knowledge, there is no large-scale benchmark for evaluating transfer-based S-HAR attacks. To fill this gap, we build the first large-scale benchmark for robust S-HAR evaluation, named RobustBenchHAR. We briefly introduce the benchmark settings here, with additional details available in Appendix C.
(A) Datasets. RobustBenchHAR incorporates three popular S-HAR datasets: NTU 60 (Shahroudy et al. 2016) , NTU 120 (Liu et al. 2019) and HDM05(Müller et al. 2007). Since the classifiers do not have the same data pre-processing setting, we unify the data format following (Wang et al. 2023). For NTU 60 and NTU 120, we subsampled frames to 60. For HDM05, we segmented the data into 60-frame samples.
(B) Evaluated Models. We evaluate TASAR in three categories of surrogate/victim models. (1) Normally trained models: We adapt 5 commonly used GCN-based models, i.e., ST-GCN (Yan, Xiong, and Lin 2018), MS-G3D (Liu et al. 2020), CTR-GCN (Chen et al. 2021), 2s-AGCN (Shi et al. 2019b), FR-HEAD (Zhou, Liu, and Wang 2023), and two latest Transformer-based models SkateFormer(Do and Kim 2024) and STTFormer(Qiu et al. 2022). To our best knowledge, this is the first work to investigate the robustness of Transformer-based S-HARs. (2) Ensemble models: an ensemble of ST-CGN, MS-G3D and DGNN (Shi et al. 2019a). (3) Defense models: We employ BEAT (Wang et al. 2023) and TRADES (Zhang et al. 2019), which all demonstrate their robustness for skeletal classifiers.
(C) Baselines. We compare with state-of-the-art (SOTA) S-HAR attacks, i.e. SMART (Wang et al. 2021) and CIASA (Liu, Akhtar, and Mian 2020). We also adopt the SOTA transfer-based attacks as baselines, including gradient-based, i.e., I-FGSM (Kurakin, Goodfellow, and Bengio 2018), MI-FGSM (Dong et al. 2018) and the latest MIG (Ma et al. 2023), input transformation method DIM (Xie et al. 2019), and ensemble-based/Bayesian attacks, i.e., ENS (Dong et al. 2018), SVRE (Xiong et al. 2022) and BA (Li et al. 2023). For a fair comparison, we ran 200 iterations for all attacks under $l_{\infty}$ norm-bounded perturbation of size 0.01. For TASAR, we use the iterative gradient attack instead of FGSM in Eq. 10.

### v5

@@ 3.3 A Post-train Bayesian Perspective on Attack
$\underset{\theta^{\prime}}{\max}\,\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)}p\left(y\mid\mathbf{x},\theta,\theta^{\prime}\right)\approx\underset{\theta_{k}^{\prime}}{\max}\,\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\mathbf{x},\theta,\theta_{k}^{\prime}\right),\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\theta),$ (3)
where $K$ is the number of appended models. Directly training such a Bayesian component is intractable, so the posterior distribution $p(\theta^{\prime}\mid\mathcal{D},\theta)$ needs to be approximated through sampling, where $p(\theta^{\prime}\mid\mathcal{D},\theta)\propto p(\mathcal{D}\mid\theta,\theta^{\prime})p(\theta^{\prime})$ and $p(\theta^{\prime})$ is the prior of appended model weights. Correspondingly, Equation 2 can be approximately solved by performing attacks on the ensemble of tiny appended models:
$\underset{\|\delta\|_{p}\leq\epsilon}{\arg\min}\,\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\tilde{\mathbf{x}},\theta,\theta_{k}^{\prime}\right),\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\theta).$ (4)
@@
3.4 Post-train Dual Bayesian Motion Attack
@@
In our preliminary experiments, we found that a naive application of post-train Bayesian attack (Equation 4) already surpassed the adversarial transfer performance of existing S-HAR attacks, which demonstrates the effectiveness of smoothening the loss surface of surrogates. However, its performance remains slightly inferior to the Bayesian attack via re-training a Bayesian surrogate (Li et al., 2023)(Equation 2). This performance gap is understandable, as we avoid the prohibitively slow process of sampling the original posterior distribution $\theta\sim p(\theta\mid\mathcal{D})$ by using a tiny Bayesian component for post-training instead. To further eliminate the trade-off between attack strength and efficiency, we propose a novel post-train dual Bayesian optimization for smoothed posterior sampling, to sample the appended models with high smoothness for better transferability (Figure 2). Moreover, unlike previous transfer-based attacks that assume each frame is independent and ignore the temporal dependency between sequences, we integrate motion dynamics information into the Bayesian attack gradient to disrupt the spatial-temporal coherence of S-HAR models. We name our method Post-train Dual Bayesian Motion Attack.
3.4.1 Post-train Dual Bayesian Optimization
@@
This motivation is based on the view that models sampled from a smooth posterior, along with the optimal approximate posterior estimating this smooth posterior, have better smoothness (Nguyen et al., 2024). To this end, we aim for proposing a smooth posterior for learning post-train BNNs, hence possibly possessing higher adversarial transferability. Specifically, inspired by the observation that randomized weights often achieve smoothed weights update (Izmailov et al., 2018; Dziugaite & Roy, 2017; Jin et al., 2023), we add Gaussian noise to smooth the appended network weights. This is achieved by a new post-train dual Bayesian optimization:
@@
$\underset{\theta^{\prime}}{\max}\,\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)}\mathbb{E}_{\Delta\theta^{\prime}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right)}p\left(y\mid\mathbf{x},\theta,\theta^{\prime}+\Delta\theta^{\prime}\right).$ (5)
@@
For any appended model sampled from the posterior, Equation 5 ensures that the neighborhood around the model parameters has uniformly low loss. We further use dual Monte Carlo sampling to approximate Equation 5:
@@
$\min_{{\theta^{\prime}_{k}}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)}\frac{1}{MK}\sum_{k=1}^{K}\sum_{m=1}^{M}L\left(\mathbf{x},y,\theta,{\theta}_{k}^{\prime}+\Delta\theta^{\prime}_{km}\right)\text{, }\Delta\theta^{\prime}_{km}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right),$ (6)
where $L$ is the classification loss. Considering dual MCMC samplings computationally intensive, we instead consider the worst-case parameters from the posterior, followed by Li et al. (2023). Hence Equation 6 can be equivalent to a min-max optimization problem, written as:
@@
$\min_{\theta^{\prime}_{k}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)}\max_{\Delta\theta^{\prime}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right)}\frac{1}{K}\sum_{k=1}^{K}L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}+\Delta\theta^{\prime}\right)\text{, }p(\Delta\theta^{\prime})\geq\xi.$ (7)
The confidence region of the Gaussian posterior is regulated by $\xi$. We discuss the sensitivity to $\xi$ in the Appendix C. The entanglement between $\theta^{\prime}$ and $\Delta\theta^{\prime}$ complicates gradient updating. To simplify this issue, we utilize Taylor expansion at $\theta^{\prime}$ to decompose the two components:
@@
$\min_{\theta^{\prime}_{k}\sim p(\theta^{\prime}\mid\mathcal{D},\theta)}\max_{\Delta\theta^{\prime}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right)}\frac{1}{K}\sum_{k=1}^{K}[L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}\right)+\nabla_{\theta^{\prime}_{k}}L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}\right)^{T}\Delta\theta^{\prime}],p(\Delta\theta^{\prime})\geq\xi.$ (8)
@@
Since $\Delta\theta^{\prime}$ is sampled from a zero-mean isotropic Gaussian distribution, the inner maximization can be solved analytically. We introduce the inference details, mathematical deduction and algorithm in Appendix B. As shown in Figure 2, the loss landscape optimized by post-train Dual Bayesian is significantly smoother than vanilla post-train Bayesian.
3.4.2 Temporal Motion Gradient in Bayesian Attack
@@
Post-train Dual Bayesian Motion Attack can be performed with gradient-based methods such as FGSM (Goodfellow et al., 2014b):
@@
$\tilde{\mathbf{x}}=\mathbf{x}+\alpha\cdot\operatorname{sign}(\sum_{k=1}^{K}\sum_{m=1}^{M}\nabla L\left(\mathbf{x},y,\theta,{\theta}_{k}^{\prime}+\Delta\theta^{\prime}_{km}\right)),$ (9)
@@
where $\alpha$ is the attack step size. Meanwhile, for notational simplicity, we notate the classification loss $L\left(\mathbf{x},y,\theta,{\theta}_{k}^{\prime}+\Delta\theta^{\prime}_{km}\right)$ as $L\left(\mathbf{x}\right)$. Assume a motion with $t$ frames $\mathbf{x}=\left[x_{1},x_{2},\cdots,x_{t}\right]$, this attack gradient consists of a set of partial derivatives over all frames $\nabla L(\mathbf{x})=\left[\frac{\partial L(\mathbf{x})}{\partial x_{1}},\frac{\partial L(\mathbf{x})}{\partial x_{2}},\cdots,\frac{\partial L(\mathbf{x})}{\partial x_{t}}\right]$. The partial derivative $\frac{\partial L(\mathbf{x})}{\partial x_{t}}$ assumes each frame is independent, ignoring the dependency between frames over time. This assumption is reasonable for attacks on static data such as PGD (Madry et al., 2017) while infeasible for skeletal motion attacks. In skeletal motion, most S-HAR models learn the spatial-temporal features (Yan et al., 2018), hence considering motion dynamics in the computing of attack gradient can disrupt the spatial-temporal coherence of these features, leading to more general transferability. To fully represent the motion dynamics, first-order (velocity) gradient $(\nabla L\left(\mathbf{x}\right))_{d1}$ and second-order (acceleration) gradient information $(\nabla L\left(\mathbf{x}\right))_{d2}$ should also be considered. To this end, we augment the original position gradient with the motion gradient, then Equation 4 becomes:
@@
$\tilde{\mathbf{x}}=\mathbf{x}+\alpha\cdot\operatorname{sign}(\sum_{k=1}^{K}\sum_{m=1}^{M}\sum_{n=0}^{2}w_{n}(\nabla L\left(\mathbf{x}\right))_{dn}),\sum_{n=0}^{2}w_{n}=1,$ (10)
@@
where $(\nabla L\left(\mathbf{x}\right))_{d0}=\nabla L(\mathbf{x})$. Motion gradient can be computed by explicit modeling (Xia et al., 2015) or implicit learning (Tang et al., 2022). Given that implicit learning requires training an additional data-driven model to learn the motion manifold, which increases computational overhead, we opt for explicit modeling. Inspired by Lu et al. (2023), we employ time-varying autoregressive models (TV-AR)(Bringmann et al., 2017) because TV-AR can effectively estimate the dynamics of skeleton sequences by modeling the temporary non-stationary signals (Xia et al., 2015). We first use first-order TV-AR($f_{d1}$) and second-order TV-AR($f_{d2}$) to model human motions respectively:
$f_{d1}:\tilde{x}^{i}_{t}$ $=A_{t}\cdot\tilde{x}^{i}_{t-1}+B_{t}+\gamma_{t},$ (11)
$f_{d2}:\tilde{x}^{i}_{t}$ $=C_{t}\cdot\tilde{x}^{i}_{t-1}+D_{t}\cdot\tilde{x}^{i}_{t-2}+E_{t}+\gamma_{t},$ (12)
@@
where the model parameters $\beta_{t}^{1}=\left[A_{t},B_{t}\right]$ and $\beta_{t}^{2}=\left[C_{t},D_{t},E_{t}\right]$ are all time-varying parameters and determined by data-fitting. $\gamma_{t}$ is a time-dependent white noise representing the dynamics of stochasticity. Using Equation 11, the first-order motion gradient can be derived as:
$\left(\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t-1}}\right)_{d1}=\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t-1}}+\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t}}\cdot A_{t}.$ (13)
@@
Similarly, second-order dynamics can be expressed as below by using Equation 12:
$\left(\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t-2}}\right)_{d2}=\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t-2}}+\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t-1}}\cdot C_{t-1}+\frac{\partial L(\tilde{\mathbf{x}}^{i})}{\partial\tilde{x}^{i}_{t}}\cdot\left(D_{t}+C_{t}\cdot C_{t-1}\right),$ (14)
@@
4 Experiments
4.1 RobustBenchHAR Settings
To our best knowledge, there is no large-scale benchmark for evaluating transfer-based S-HAR attacks. To fill this gap, we build the first large-scale benchmark for robust S-HAR evaluation, named RobustBenchHAR. We briefly introduce the benchmark settings here, with additional details available in Appendix D.
(A) Datasets. RobustBenchHAR incorporates three popular S-HAR datasets: NTU 60 (Shahroudy et al., 2016) , NTU 120 (Liu et al., 2019) and HDM05(Müller et al., 2007). Since the classifiers do not have the same data pre-processing setting, we unify the data format following (Wang et al., 2023). For NTU 60 and NTU 120, we subsampled frames to 60. For HDM05, we segmented the data into 60-frame samples.
(B) Evaluated Models. We evaluate TASAR in three categories of surrogate/victim models. (1) Normally trained models: We adapt 5 commonly used GCN-based models, i.e., ST-GCN (Yan et al., 2018), MS-G3D (Liu et al., 2020b), CTR-GCN (Chen et al., 2021), 2s-AGCN (Shi et al., 2019a), FR-HEAD (Zhou et al., 2023), and two latest Transformer-based models SkateFormer(Do & Kim, 2024) and STTFormer(Qiu et al., 2022). To our best knowledge, this is the first work to investigate the robustness of Transformer-based S-HARs. (2) Ensemble models: an ensemble of ST-CGN, MS-G3D and DGNN (Shi et al., 2019b). (3) Defense models: We employ BEAT (Wang et al., 2023) and TRADES (Zhang et al., 2019a), which all demonstrate their robustness for skeletal classifiers.
(C) Baselines. We compare with state-of-the-art (SOTA) S-HAR attacks, i.e. SMART (Wang et al., 2021) and CIASA (Liu et al., 2020a). We also adopt the SOTA transfer-based attacks as baselines, including gradient-based, i.e., I-FGSM (Kurakin et al., 2018), MI-FGSM (Dong et al., 2018) and the latest MIG (Ma et al., 2023), input transformation method DIM (Xie et al., 2019), and ensemble-based/Bayesian attacks, i.e., ENS (Dong et al., 2018), SVRE (Xiong et al., 2022) and BA (Li et al., 2023). For a fair comparison, we ran 200 iterations for all attacks under $l_{\infty}$ norm-bounded perturbation of size 0.01. For TASAR, we use the iterative gradient attack instead of FGSM in Equation 10.
Table 1: The attack success rate(%) of untargeted transfer-based attacks on NTU60 and NTU120. ’Ave’ was calculated as the average transfer success rate over all target models except for the surrogate.’SFormer’ represents SkateFormer and MI stands for MI-FGSM.